OpenStudy (kimberly_pr):

What is the cube root of 1,600?

OpenStudy (braydonlevi99):

What can you multiply three times in order to get 1,600?

OpenStudy (mathstudent55):

Do you need an exact answer with a simplified radical or an approximation?

OpenStudy (kimberly_pr):

Well it's either 40, 8 root 5, 4 times cube root 5, or 4 times cube root 25

jimthompson5910 (jim_thompson5910):

Hints: \[\Large \sqrt[3]{1600}=\sqrt[3]{64*25}\] \[\Large 64 = 4*4*4 = 4^3\]

OpenStudy (mathstudent55):